Ordered: Acetaminophen 15mg/kg/dose every 4-6hrs as needed for fever. Available: Acetaminophen 160/5mL. The patient weighs 70 lbs. How many mL will you administer for one dose? ( Round to the nearest tenth)
Answer #1

Ordered: Acetaminophen 15mg/kg/dose every 4-6hrs as needed for fever

The patient weighs 70 lbs= 70/2.2 kg= 31.81 kg (since 1 kg = 2.2 lbs or pounds)

Hence as per patient's weight required dose= 31.81X 15= 477.15‬ mg/ dose every 4-6hrs

Available: Acetaminophen 160/5mL.

Means 32 mg/ml

How many mL will you administer for one dose?

32 mg is available in 1 ml

1 mg will be available in 1/32 ml

Hence 477.15‬ mg will be available in = (1/32) X 477.15 ml

= 14.91 ml= 15 ml (as rounded to nearest tenth)